# Towards Generalizing to Unseen Domains with Few Labels

### v1

@@ Abstract
We approach the challenge of addressing semi-supervised domain generalization (SSDG). Specifically, our aim is to obtain a model that learns domain-generalizable features by leveraging a limited subset of labelled data alongside a substantially larger pool of unlabeled data. Existing domain generalization (DG) methods which are unable to exploit unlabeled data perform poorly compared to semi-supervised learning (SSL) methods under SSDG setting. Nevertheless, SSL methods have considerable room for performance improvement when compared to fully-supervised DG training. To tackle this underexplored, yet highly practical problem of SSDG, we make the following core contributions. First, we propose a feature-based conformity technique that matches the posterior distributions from the feature space with the pseudo-label from the model’s output space. Second, we develop a semantics alignment loss to learn semantically-compatible representations by regularizing the semantic structure in the feature space. Our method is plug-and-play and can be readily integrated with different SSL-based SSDG baselines without introducing any additional parameters. Extensive experimental results across five challenging DG benchmarks with four strong SSL baselines suggest that our method provides consistent and notable gains in two different SSDG settings. Our code will is available at FBC-SA.
11footnotetext: Mohamed bin Zayed University of Artificial Intelligence
@@ 3.1 Preliminaries
Problem Settings: Similar to multi-source DG settings, let us characterize each domain $d$ by $d=\{(\mathbf{x}_{i}^{d},y_{i}^{d})\}_{i=1}^{n}$, where $\mathbf{x}^{d}_{i}\in\mathbb{R}^{C\times H\times W}$ is an input image, $y^{d}_{i}$ is the corresponding label, and comprised of $n$ [^1] The value of $n$ can be different for each domain. independent and identically distributed (i.i.d) examples drawn from a joint probability distribution $\mathcal{P}(\mathcal{X}^{d},\mathcal{Y}^{d})$ for all possible training (source) domains $d\in\{d_{1},...,d_{tr}\}$. $\mathcal{X}^{d}$ is an input space over which the domain $d$ is defined and $\mathcal{Y}^{d}$ is the corresponding label space. Here, we consider the distribution shift in $\mathcal{P}(\mathcal{X}^{d})$ while $\mathcal{P}(\mathcal{Y}^{d})$ shares the same label space $\mathcal{Y}$ where $y\in\mathcal{Y}=\{1,2,..,C\}$ is an associated class label.
@@ 4 Experiments
Datasets, training and implementation details: We utilize PACS [23], OfficeHome [38], Digits [58], TerraIncognita [5] and VLCS [36] datasets which are widely used to report domain generalization performance. For a detailed description of the datasets, refer to the supplementary materials. We conduct experiments under two settings; 10 labels and 5 labels per class while labeled images are selected randomly. The latter setting is more challenging due to the extreme scarcity of labeled data. Following [61], we randomly sample 16 labeled and 16 unlabeled images from each source domain to construct a minibatch. The labeled subset of minibatch is used to calculate the supervised loss while the (complete) minibatch, including both the labeled images (with ground truth labels dropped) and unlabeled images, are utilized to calculate all the unsupervised losses [61]. We use ImageNet[12] pre-trained ResNet-18[18] as the backbone architecture and a single-layer MLP head as the classifier. We use SGD as the optimizer with an initial learning rate of 0.003 for the backbone and 0.01 for the classifier, respectively. Both learning rates are decayed using the cosine annealing rule and we train all the models for 20 epochs on all datasets except TerraIncognita (trained for 10 epochs). We set $n=\lceil\frac{|C|}{2}\rceil$, from sec. 3.3, in all experiments.
@@
Baselines: Since our method combines SSL with DG, we select state-of-the-art methods in both paradigms. We choose EntMin [16], MeanTeacher [35], FixMatch [33], FlexMatch [52] and FreeMatch [43] methods in SSL. StyleMatch[61] is selected as an SSDG baseline as it shows promising performance under 5 and 10 labels settings. Furthermore, we chose ERM[37] in DG as it shows competitive performance against many existing DG methods [17]. For a detailed comparison with existing DG methods, we refer to the supplementary materials.
@@ 4.1 Results
It should be noted that StyleMatch is unable to improve on its baseline FixMatch when there are no style distributions present in the source domains (on datasets such as Digits and TerraIncognita) where our proposed method demonstrates significant gains with over 4.0% improvement in accuracy. As our approach is model agnostic, it can be seamlessly integrated with different SSL and SSDG baselines. Further, the improvements over the baselines are consistent across all the DG datasets under both 10 labels and 5 labels per class settings.
@@
Table 3: Performance contribution of components in our approach on PACS (5 labels per class).
@@ 5 Conclusion and Limitations
We approach the relatively unexplored problem of semi-supervised domain generalization (SSDG) and propose a new method, built with feature-based conformity and semantics alignment constraint modules, towards addressing the key challenges in SSDG. The feature-based conformity mechanism aligns the posterior distributions from two views, while the semantics alignment constraint further boosts the effectiveness of feature-based conformity by regularizing the semantic layout of feature space. Our approach is plug-and-play, parameter-free and model-agnostic, so it can be seamlessly integrated into different baselines as validated in our results. Extensive experiments on different challenging DG benchmarks suggest that our method is capable of delivering a consistent and notable gain over four recent baselines. An aspect to consider is that our approach may necessitate significant adjustments to accommodate semi-supervised single-source DG.
Supplementary Material
6 Datasets Description
PACS: PACS dataset consists of four domains: photo (1,670 images), art painting (2,048 images), cartoon (2,344 images) and sketch (3,929 images), and seven classes: dog, elephant, giraffe, guitar, horse, house, and person.
OfficeHome: The OfficeHome dataset also consists of four domains: art, clipart, product, and the real world with 65 different classes whose total number of images adds up to around 15,500.
Digits-DG: The Digits-DG dataset consists of four domains: MNIST, MNIST-M, SVHN, and SYN, and 10 classes from 0 to 9. MNIST comprises hand-written digits, MNIST-M is a variant of MNIST with a blended background, SVHN has street-view house numbers and SYN consists of synthetic digits.
VLCS: VLCS dataset consists of four domains: Caltech, Labelme, Pascal, and Sun with 5 classes: bird, car, chair, dog, and person
Terra Incognita: This dataset contains images belonging to four domains collected from different locations namely locations 38, 43,46 and 100 with 10 classes: bird, bobcat, cat, coyote, dog, empty, opossum, rabbit, raccoon, and squirrel.
Figure 6 displays randomly selected samples from each dataset.
(a) PACS
(b) OfficeHome
(c) VLCS
(d) Digits-DG
(e) Terra Incognita
Figure 6: Example images from different DG datasets used in our experiments.
7 Comparison with DG baselines
We choose ERM [37], MixUp [54], GroupDRO [29], CrossGrad [31], DAELDG [60] , DDAIG [58] and DomainMix [34] as DG baselines to compare our method on PACS, OfficeHome, Digits-DG, VLCS and TerraIncognita datasets under both 5 labels (see Tab. 5) and 10 labels (see Tab. 6) settings.
8 Additional class similarity matrices
To show the effectiveness of our algorithm, we show cosine similarity between the class means on PACS dataset in the main manuscript. In this supplementary, we also show this on Digits-DG dataset (see Fig. 7). We see that, our losses encourage orthogonality among features with different class labels, and hence better discrimination in the feature space under different domain shifts and limited labels. Especially in Mnist-m, and Syn domains the similarities are notably higher compared to the Fixmatch baseline.
Model PACS OH VLCS DigitsDG TerraInc.
ERM $51.2\pm 3.0$ $51.7\pm 0.6$ $67.2\pm 1.8$ $22.7\pm 1.0$ $22.9\pm 3.0$
MixUp $45.3\pm 3.8$ 52.7 $\pm$0.6 $69.9\pm$ 1.3 $21.7\pm$ 1.9 21.0$\pm$2.9
GroupDRO $48.2\pm 3.6$ 53.8±$\pm 0.6$ $69.8\pm 1.2$ $23.1\pm$1 9 $22.4\pm 3.1$
CrossGrad $50.6\pm 3.4$ $51.6\pm 0.9$ $68.1\pm 1.6$ $22.8\pm 0.4$ $21.4\pm 2.3$
DAELDG $42.7\pm 2.7$ $47.3\pm 0.6$ $61.7\pm 1.9$ $22.3\pm 1.0$ $25.0\pm 3.0$
DDAIG $50.5\pm 3.0$ $50.6\pm 0.7$ $65.2\pm 2.2$ $23.2\pm 1.8$ $31.7\pm 3.1$
DomainMix $46.3\pm 3.5$ $49.9\pm 0.6$ $68.3\pm 0.7$ $20.6\pm 1.4$ $22.9\pm 0.9$
Table 5: DG accuracy (%) under SSDG settings (5 labels per class). Average over 5 independent seeds is reported.
Model PACS OH VLCS DigitsDG TerraInc.
ERM $59.8\pm 2.3$ $56.7\pm 0.8$ $68.0\pm 0.3$ $29.1\pm 2.9$ $23.5\pm 1.2$
MixUp $58.5\pm 2.2$ $57.2\pm 0.6$ $69.6\pm 1.0$ $29.7\pm 3.1$ $24.8\pm 3.3$
GroupDRO $57.3\pm 1.2$ $57.8\pm 0.4$ $69.4\pm 0.9$ $31.5\pm 2.5$ $25.8\pm 3.3$
CrossGrad $59.7\pm 1.5$ $56.7\pm 0.4$ $67.9\pm 0.6$ $30.3\pm 2.7$ $22.6\pm 0.9$
DAELDG $53.7\pm 2.1$ $54.8\pm 0.3$ $68.3\pm 1.3$ $28.6\pm 1.5$ $25.5\pm 2.6$
DDAIG $59.6\pm 1.6$ $55.1\pm 0.1$ $68.5\pm 1.0$ $29.4\pm 3.0$ $23.5\pm 3.0$
DomainMix $58.0\pm 1.9$ $55.4\pm 0.4$ $69.5\pm 0.8$ $24.6\pm 1.1$ $23.3\pm 2.1$
Table 6: DG accuracy (%) under SSDG settings (10 labels per class). Average over 5 independent seeds is reported.
Figure 7: Comparison of cosine similarity of mean class features between Fixmatch and our method on Digits-DG.
9 Confusion Matrices
We plot the confusion matrices comparison between the Fixmatch baseline and our method on PACS (Fig. 8), and Digits-DG (Fig. 9) datasets. Compared to FixMatch, our approach shows improved class-wise accuracy in both datasets.
Figure 8: Confusion matrix comparison between Fixmatch and our method on PACS.
Figure 9: Confusion matrix for Fixmatch and our method on Digits-DG.
10 Additional Feature Visualizations
In addition to the feature visualization on PACS dataset (main manuscript), we visualize features on Digits-DG dataset in Fig. 10 for Fixmatch and our method. In our approach, we observe that intra-class features are closer and inter-class features are far apart. We note that, compared to the baseline, the classes are well-separated in our method. Also, the comparison of cosine similarity of mean class features in Fig. 7 validates this further.
11 Comparison with other backbones
For a fair comparison, we used the same backbone (ResNet18) as used in StyleMatch. To evaluate our method further, we analyze the performance OfficeHome dataset with several stronger backbones such as ImageNet pre-trained ResNet50, ResNet101, ViT-S/32, ViT-B/32, and CLIP pre-trained ViT-B/32 (CLIP-B/32) in Tab. 7. Our method consistently outperforms the baseline even with other stronger backbones.
12 Scaling of performance with # labels
Tab. 8 provides results with increasing number of per-class labels. We see that the performance of our method improves upon increasing the number of per-class labels and in all per-class labels settings, it is higher than the other methods. Furthermore, with just 100 per-class labels, our method performance is better than the fully supervised ERM (i.e. with all labels) that obtains 80.0±${}_{\pm 0.5}$.
13 Runtime and Memory overhead comparison
To evaluate the effectiveness of our method, we compare the runtime and memory overhead (Tab. 9.) Our method adds a little runtime overhead (20%) over the baseline FixMatch compared to the existing SSDG method, StyleMatch (117%) which has the same baseline (FixMatch).
Algorithm RN 50 RN 101 Vit-S/32 Vit-B/32 CLIP-B/32
FixMatch[32] 61.3±${}_{\pm 0.4}$ 62.8±${}_{\pm 0.2}$ 63.7±${}_{\pm 0.5}$ 72.0±${}_{\pm 0.4}$ 75.3±${}_{\pm 0.6}$
FixM. +Ours 62.1±${}_{\pm 0.4}$ 64.2±${}_{\pm 0.1}$ 64.4±${}_{\pm 0.3}$ 72.9±${}_{\pm 0.3}$ 78.9±${}_{\pm 0.4}$
Table 7: Results with different backbones on the Office Home dataset (10 labels per class)
Algorithm 5 10 25 50 100
ERM[35] 51.2±${}_{\pm 1.0}$ 59.8±${}_{\pm 2.5}$ 66.7±${}_{\pm 2.2}$ 71.2±${}_{\pm 1.9}$ 75.7±${}_{\pm 1.6}$
FixMatch[32] 72.8±${}_{\pm 1.2}$ 76.6±${}_{\pm 1.2}$ 77.6±${}_{\pm 1.4}$ 78.7±${}_{\pm 0.4}$ 79.4±${}_{\pm 1.4}$
FixM.+Ours 77.3±${}_{\pm 1.1}$ 78.2±${}_{\pm 1.2}$ 79.3±${}_{\pm 1.8}$ 79.6±${}_{\pm 1.0}$ 80.4±${}_{\pm 0.6}$
Table 8: Results with different per class labels on PACS
Algorithm s/epoch Overhead GPU Mem(MB) Overhead
FixMatch[32] 44.4 - 5595 -
StyleMatch[57] 96.38 +117.07 % 7561 +35.13 %
FixM.+Ours 52.85 +20.11 % 7647 +36.67 %
Table 9: Runtime (s/epoch) and memory (MB) overhead
Figure 10: Feature visualization using tSNE for Fixmatch and ours on Digits-DG. Our method facilitates learning more discriminative features under various domain shifts and limited labels.

### v3

@@ Abstract
We approach the challenge of addressing semi-supervised domain generalization (SSDG). Specifically, our aim is to obtain a model that learns domain-generalizable features by leveraging a limited subset of labelled data alongside a substantially larger pool of unlabeled data. Existing domain generalization (DG) methods which are unable to exploit unlabeled data perform poorly compared to semi-supervised learning (SSL) methods under SSDG setting. Nevertheless, SSL methods have considerable room for performance improvement when compared to fully-supervised DG training. To tackle this underexplored, yet highly practical problem of SSDG, we make the following core contributions. First, we propose a feature-based conformity technique that matches the posterior distributions from the feature space with the pseudo-label from the model’s output space. Second, we develop a semantics alignment loss to learn semantically-compatible representations by regularizing the semantic structure in the feature space. Our method is plug-and-play and can be readily integrated with different SSL-based SSDG baselines without introducing any additional parameters. Extensive experimental results across five challenging DG benchmarks with four strong SSL baselines suggest that our method provides consistent and notable gains in two different SSDG settings. Our code is available at FBC-SA.
**footnotetext: Equally contributing authors
@@ 3.1 Preliminaries
Problem Settings: We adapt some notation from [17]. Similar to multi-source DG settings, let us characterize each domain $d$ by $d=\{(\mathbf{x}_{i}^{d},y_{i}^{d})\}_{i=1}^{n}$, where $\mathbf{x}^{d}_{i}\in\mathbb{R}^{C\times H\times W}$ is an input image, $y^{d}_{i}$ is the corresponding label, and comprised of $n$***The value of $n$ can be different for each domain. independent and identically distributed (i.i.d) examples drawn from a joint probability distribution $\mathcal{P}(\mathcal{X}^{d},\mathcal{Y}^{d})$ for all possible training (source) domains $d\in\{d_{1},...,d_{tr}\}$. $\mathcal{X}^{d}$ is an input space over which the domain $d$ is defined and $\mathcal{Y}^{d}$ is the corresponding label space. Here, we consider the distribution shift in $\mathcal{P}(\mathcal{X}^{d})$ while $\mathcal{P}(\mathcal{Y}^{d})$ shares the same label space $\mathcal{Y}$ where $y\in\mathcal{Y}=\{1,2,..,C\}$ is an associated class label.
@@ 4 Experiments
Datasets, training and implementation details: We utilize PACS [23], OfficeHome [36], Digits [55], TerraIncognita [5] and VLCS [34] datasets which are widely used to report DG performance. For a detailed description of the datasets, refer to the supplementary materials. We conduct experiments under two settings; 10 labels and 5 labels per class while labeled images are selected randomly. The latter setting is more challenging due to the extreme scarcity of labeled data. Following [57], we randomly sample 16 labeled and 16 unlabeled images from each source domain to construct a minibatch. The labeled subset of minibatch is used to calculate the supervised loss while the (complete) minibatch, including both the labeled images (with ground truth labels dropped) and unlabeled images, are utilized to calculate all the unsupervised losses [57]. We use ImageNet[12] pre-trained ResNet-18[18] as the backbone architecture and a single-layer MLP head as the classifier. We use SGD as the optimizer with an initial learning rate of 0.003 for the backbone and 0.01 for the classifier, respectively. Both learning rates are decayed using the cosine annealing rule and we train all the models for 20 epochs on all datasets except TerraIncognita (trained for 10 epochs). We set $n=\lceil\frac{|C|}{2}\rceil$, from sec. 3.3, in all experiments.
@@
Baselines: Since our method combines SSL with DG, we select state-of-the-art methods in both paradigms. We choose EntMin [16], MeanTeacher [33], FixMatch [32], FlexMatch [50] and FreeMatch [41] methods in SSL. StyleMatch[57] is selected as an SSDG baseline as it shows promising performance under 5 and 10 labels settings. Further, we chose ERM[35] in DG as it shows competitive performance against many existing DG methods [17]. For a detailed comparison with existing DG methods, see supplementary materials.
@@ 4.1 Results
Note that, StyleMatch is unable to improve on its baseline FixMatch when there are no style distributions present in the source domains (on datasets such as Digits and TerraIncognita) where our proposed method demonstrates significant gains with over 4.0% improvement. As our approach is model agnostic, it can be seamlessly integrated with different SSL and SSDG baselines. Further, the improvements over the baselines are consistent across all the DG datasets under both 10 and 5 labels per class settings.
@@
Table 3: Ablation study on PACS (5 labels per class).
@@ 5 Conclusion and Limitations
We approach the relatively unexplored problem of semi-supervised domain generalization (SSDG) and propose a new method, built with feature-based conformity and semantics alignment constraint modules, towards addressing the key challenges in SSDG. The feature-based conformity mechanism aligns the posterior distributions from two views, while the semantics alignment constraint further boosts the effectiveness of feature-based conformity by regularizing the semantic layout of feature space. Our approach is plug-and-play, parameter-free and model-agnostic, so it can be seamlessly integrated into different baselines as validated in our results. Extensive experiments on different challenging DG benchmarks show that our method delivers a consistent and notable gain over four recent baselines. An aspect to consider is that our approach may necessitate significant adjustments to accommodate semi-supervised single-source DG.